Find the time required for an investment of 5000 dollars to grow to 7200 dollars at an interest rate of 7.5 percent per year, compounded quarterly.
Round your answer to two decimal places
Answer #1

We use the formula:
A=P(1+r/4)^4n
where
A=future value
P=present value
r=rate of interest
n=time period.

7200=5000*(1+0.075/4)^4n

(7200/5000)=(1+0.075/4)^4n

1.44=(1.01875)^4n

Taking log on both sides;

log 1.44=4n*log (1.01875)

n=1/4[log 1.44/log (1.01875)]

=4.91 years(Approx).